OpenStudy (anonymous):

A truck travels 400 miles on level road in the same amount of time it travels 160 miles on a mountainous road. If the rate of the truck is 30mile per hour less on the mountain road than on the level road, find its rate on both roads.

OpenStudy (wolf1728):

R1 = Rate1 R2 = Rate2 The time is A) 400/R1 = 160/R2 Rates B) R1 = R2 +30 Putting R2 +30 into equation A 400/(R2+30) = 160/R2 400R2 = 160R2 +4800 240R2 = 4800 R2 = 20 R1 = 50 Time 400/50 = 8 Hours 160/20 = 8 Hours

OpenStudy (anonymous):

thank you